# On the Weight Spectrum of Rate-Compatible Polar Codes

### v2

@@ Abstract
The weight spectrum plays a crucial role in the performance of error-correcting codes. Despite substantial theoretical exploration into polar codes with mother code length, a framework for the weight spectrum of rate-compatible polar codes remains elusive. In this paper, we address this gap by enumerating the number of minimum-weight codewords for quasi-uniform punctured, Wang-Liu shortened, and bit-reversal shortened polar codes. Additionally, we propose efficient algorithms for computing the average spectrum of random upper-triangular pre-transformed shortened and punctured polar codes. Notably, our algorithms operate with polynomial complexity relative to the code length. Simulation results affirm that our findings yield a precise estimation on the performance of rate-compatible polar codes.
@@ I Introduction
The weight spectrum has a significant impact on the ML decoding performance, which can be approximated through the union bound based on the number of low-weight codewords [14]. However, in the general case, the complexity of computing the exact weight spectrum grows exponentially with code length. Various efforts are made to analyse the weight spectrum of polar codes. The authors in [15] regarded polar codes as decreasing monomial codes and applied lower triangular affine (LTA) automorphisms to determine the number of codewords with the minimum weight $w_{\min}$. Building upon this foundation, recent studies quantified the number of codewords with weights $1.5w_{\min}$ [16] and those with weights below $2w_{\min}$ [17].
@@
Algorithmic approaches for approximating the weight spectrum of polar codes are developed. For instance, [18] proposed to employ SCL decoding with a large list size at high signal-to-noise ratios (SNR) to collect low-weight codewords. Subsequently, [19] refined this method by optimizing memory usage. In parallel, probabilistic approaches with polynomial complexity were introduced in [20, 21] to approximate the weight spectrum of polar codes to strike a balance between accuracy and computational efficiency.
@@
To improve the weight spectrum of polar codes, pre-transformed polar codes [22], including CRC-Aided (CA) polar codes[2], parity-check (PC) polar codes[23] and polarization-adjusted convolutional (PAC) codes[24] were proposed. It has been proved that such upper-triangular pre-transformations do not decrease the code distance [22]. The number of minimum-weight codewords for pre-transformed polar codes was calculated in [25]. Furthermore, the authors in [26] presented a simplified algorithm to calculate the exact weight spectrum of both original and specific pre-transformed polar codes using the coset method with an exponential complexity. The average weight spectrum of CA-polar codes, based on the input-output weight enumerating function of original polar codes, was researched in [27]. However, determining the input-output weight enumerating function for original polar codes remains challenging. In [28, 29], the authors proposed efficient recursive formulas to calculate the average weight spectrum of pre-transformed polar codes with polynomial complexity.
@@
Moreover, a method for computing the weight spectrum of rate-compatible polar codes was proposed in [30]. The authors introduced a technique to enumerate all codewords up to a certain Hamming weight for general binary linear block codes. The complexity of using this method is determined by the number of low-weight codewords in the component codes. Although the complexity for polar codes is lower than that of general linear codes, it still exceeds polynomial of code length.
@@
In this paper, we compute the number of low-weight codewords for QUP, Wang-Liu shortened, and bit-reversal shortened polar codes. We also provide iterative formulas for calculating the average weight spectrum of random pre-transformed rate-compatible polar codes. The algorithms operate with polynomial complexity with respect to code length. This computational characteristic ensures that our methods remain scalable and practical for use with polar codes of varying lengths.
@@ II-A Polar codes as monomial codes
Let $f=x_{1}^{a_{1}}...x_{m}^{a_{m}}$ be a monomial in $\mathcal{M}$, the degree of $f$ is defined as $\text{deg}(f)$, i.e., the Hamming weight of $(a_{1},...,a_{m})$. In particular, 1 is a monomial with degree zero. We say monomial $f$ is a factor of monomial $g$ if $g=fh$.
@@
and the degree of the polynomial $p=\sum_{e\in\mathcal{M}}u_{f}f$ is defined as $\text{deg}(p)=\max_{a_{f}\neq 0}\{\text{deg}(f)\}$. A polynomial $p$ is said to be linear if deg$(p)=1$.
@@
where the element $p(\bm{a})$ is placed at the $D(\bm{a})$-$th$ position in $\text{ev}(p)$.
Denote $f_{i}=x_{1}^{a_{1}}...x_{m}^{a_{m}}$ with $i=D(a_{1},...,a_{m})$. Then $\text{ev}(f_{i})$ is exactly the $i$-$th$ row of $\bm{F}_{N}$. Therefore, for a codeword $\bm{c}=\bm{u}\bm{F}_{N}\in C(\mathcal{I})$, we have $\bm{c}=\text{ev}(p)$ where $p=\sum_{i=1}^{N}u_{i}f_{i}$. In this paper, we use $p$ to represent codewords $\bm{c}=\text{ev}(p)$. The Hamming weight of $p$ is $\text{wt}(p)\triangleq\text{wt}(\bm{c})$, and $p_{[a,b]}\triangleq\bm{c}_{a}^{b}=(c_{a},c_{a+1},\dots,c_{b})$.
@@
If the maximum degree of monomials in $\mathcal{I}$ is $r$, we call the monomial code $C(\mathcal{I})$ is $r$-$th$ order.
@@
If the information set is =$\mathcal{I}=\{1,x_{1},x_{2},x_{3}\}$, then the generator matrix of the polar code $C(\mathcal{I})$ is
@@
The polynomial $p=x_{3}+x_{1}+1$ is a codeword in $(C(\mathcal{I}))$, which represents the codeword $\text{ev}(p)=(1,0,1,0,1,0,0,1)$.
@@ II-E Automorphism group of decreasing monomial codes
The automorphism group Aut$(C)$ is the subgroup of $S_{N}$ containing all the automorphisms of $C$.
@@
A matrix $\bm{A}$ is lower-triangular if $A_{i,i}=1$ and $A_{i,j}=0$ for all $j>i$. The group LTA$(m)$ consists of all the affine transformations $({\bm{A}},{\bm{b}})$ where $\bm{A}$ is lower-triangular. Similarly, a matrix $\bm{A}$ is upper-triangular if $A_{i,i}=1$ and $A_{i,j}=0$ for all $j<i$.
Let $\bm{s}=(s_{1},\dots,s_{l})$ be a positive integer vector with $s_{1}+\dots+s_{l}=m$. BLTA$(\bm{s})$ is a BLTA group consisting of all the affine transformations $({\bm{A}},{\bm{b}})$ where ${\bm{A}}$ can be written as a block matrix of the following form
@@
In [15], it is proved that LTA$(m)$ belongs to the automorphism group of a decreasing monomial code. Later, a larger BLTA group is known to be the affine automorphism group of any specific decreasing monomial code in [34, 33].
@@
When $m=3$, let $\bm{A}=\bm{I}_{3}$ be the identity matrix and $\bm{b}=(1,1,1)$. Then $(\bm{A},\bm{b})$ transforms $x_{i}$ to $x_{i}+1$, which means it interchanges the $i=D((x_{1},x_{2},x_{3}))$-$th$ bit and the $(N+1-i)=D((x_{1}+1,x_{2}+1,x_{3}+1))$-$th$ bit for all $1\leq i\leq N$. The corresponding permutation is $(8,7,6,5,4,3,2,1)$.
@@ II-F Puncturing and shortening
Denote $[N]=\{1,2,\cdots,N\}$, the difference set of two sets $X$ and $Y$ is denoted as $X/Y$, i.e., $X/Y=\{i\in X\mid i\notin Y\}$. For a codeword $\bm{c}\in\mathbb{F}_{2}^{N}$ and set $X\subseteq[N]$, $\bm{c}_{X}$ is the codeword $\bm{c}$ restricted on $X$, i.e., $\bm{c}_{X}=(c_{i})_{i\in X}$.
@@
Denote the quasi-uniform (QU) sequence to be $\bm{q}=(1,2,\dots,N)$. The QUP pattern $X_{i}$ [4] is the set of the first $i$ bits in $\bm{q}$, i.e., $\{1,2,\dots,i\}$. The QUP polar code with length $N-i$ and information set $\mathcal{I}$ is denoted by $C_{P}(\mathcal{I},X_{i})$. The Wang-Liu shortening pattern $Y_{i}$ [5] is the set of the last $i$ bits in $\bm{q}$, i.e., $\{N-i+1,N-i+2,\dots,N\}$. The Wang-Liu shortened polar code with length $N-i$ and information set $\mathcal{I}$ is denoted as $C_{S}(\mathcal{I},Y_{i})$.
Denote the bit-reversal sequence $\bm{q}^{\prime}$ to be the bit reversal of $\bm{q}$, i.e., $\bm{q}^{\prime}_{D(a_{m},\dots,a_{1})}=\bm{q}_{D(a_{1},\dots,a_{m})}=D(a_{1},\dots,a_{m})$. The bit-reversal shortening pattern $Y^{\prime}_{i}$ is the last $i$ bits of $\bm{q}^{\prime}$, and the bit-reversal shortened polar code with length $N-i$ and information set $\mathcal{I}$ is denoted as $C_{S}(\mathcal{I},Y_{i}^{\prime})$.
@@
Note that when the puncturing pattern $X$ complies with binary domination [35], the symmetric capacities of monomial $x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ with $D(a_{1},\dots,a_{m})\in X$ are degraded to zero. When the shortening pattern $Y$ complies with binary domination, the coefficient of monomial $x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ with $D(a_{1},\dots,a_{m})\in Y$ are frozen to zero. Since QUP, Wang-Liu shortening and bit-reversal shortening patterns all comply with binary domination, when $X$ (resp. $Y$) is the puncturing (resp. shortening) pattern, the coefficient of monomial $x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ with $D(a_{1},\dots,a_{m})\in X$ (resp. $Y$) in $p$ is zero. Therefore, with little abuse of notation, when the $D(a_{1},\dots,a_{m})$-$th$ bit is punctured (resp. shortened), we also say the monomial $x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ is punctured (resp. shortened).
In this paper, we focus on the minimum-weight codewords of decreasing QUP, Wang-Liu shortened and bit-reversal shortened polar codes. We say a shortened polar code is decreasing if the union of the information set and the shortened set is decreasing.
@@
In section III and section IV, we calculate the number of minimum-weight codewords for QUP, Wang-Liu shortened and bit-reversal shortened polar codes. We provide an outline of the proof in this subsection. Define $T_{P}(f,X)=\{p_{[N]/X}\mid p\in T(f)\}$, i.e., the set of punctured codewords in $T(f)$ with the puncturing pattern $X$, and $T_{S}(f,Y)=\{p_{[N]/Y}\mid p\in T(f),p_{Y}=\bm{0}\}$, i.e., the set of codewords in $T(f)$ with the shortening pattern $Y$. For $f=x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$, we only need to study $T_{P}(f,X)$ (resp. $T_{P}(f,Y)$) for $D(a_{1},\dots,a_{m})\notin X$ (resp. $Y$) due to Remark 1.
In section III, we calculate the number of minimum-weight codewords for bit-reversal shortened polar codes $C_{S}(\mathcal{I},Y^{\prime}_{i})$. $T_{S}(f,Y^{\prime}_{i})$ is a subset of $T(f)$ with the coefficients of monomials in $Y^{\prime}_{i}$ being 0. As shortening does not change the codeword weight, the set of minimum-weight codewords in $C_{S}(\mathcal{I},Y^{\prime}_{i})$ is $\bigcup_{f\in\mathcal{I}_{r}}T_{S}(f,Y^{\prime}_{i})$. We demonstrate that $T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})$ are not empty if and only if the $i$-$th$ shortened monomial is a factor of $f$. Moreover, at each time there are $|T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})|=\lambda_{f}/2^{r}$ polynomials removed. As $|T_{S}(f,\varnothing)|=|T(f)|=\lambda_{f}$, $|T_{S}(f,Y^{\prime}_{i})|$ can be calculated recursively by the number of shortened bits.
In section IV, we analyse the number of minimum-weight codewords for QUP and Wang-Liu shortened polar codes. For QUP, define $M_{P}(\mathcal{I},X_{i})$ to be the set of minimum-weight codewords in $C_{P}(\mathcal{I},X_{i})$. We prove that a codeword in $M_{P}(\mathcal{I},X_{i})$ must also be a minimum-weight codeword before puncturing, i.e., belongs to $\bigcup_{f\in\mathcal{I}}T_{P}(f,X_{i})$. Therefore, we compute the weight spectrum of $\bigcup_{f\in\mathcal{I}}T_{P}(f,X_{i})$ iteratively from the weight spectrum of $T_{P}(f^{\prime},X_{j})$ where $f^{\prime}\preccurlyeq f$ and $j<i$. Additionally, the number of minimum-weight codewords in Wang-Liu shortened polar codes can be inferred from that in QUP code by symmetry.

### v3

@@ Abstract
The weight spectrum plays a crucial role in the performance of error-correcting codes. Despite substantial theoretical exploration of polar codes with mother code length, a framework for the weight spectrum of rate-compatible polar codes remains elusive. In this paper, we address this gap by presenting the theoretical results for enumerating the number of minimum-weight codewords for quasi-uniform punctured, Wang-Liu shortened, and bit-reversal shortened decreasing polar codes. Additionally, we propose efficient algorithms for computing the average spectrum of random upper-triangular pre-transformed shortened and punctured polar codes. Notably, our algorithms operate with polynomial complexity relative to the code length. Simulation results affirm that our findings yield a precise estimation of the performance of rate-compatible polar codes.
@@ I Introduction
The weight spectrum has a significant impact on the ML decoding performance, which can be approximated through the union bound based on the number of low-weight codewords [14]. However, in the general case, the complexity of computing the exact weight spectrum grows exponentially with the code length. Various efforts have been made to analyse the weight spectrum of polar codes. The authors in [15] regarded polar codes as decreasing monomial codes and applied lower triangular affine (LTA) automorphisms to determine the number of codewords with the minimum weight $w_{\min}$. Building upon this foundation, recent studies quantified the number of codewords with weights of $1.5w_{\min}$ [16] and those with weights below $2w_{\min}$ [17].
@@
Algorithmic approaches for approximating the weight spectrum of polar codes are developed. For instance, [18] proposed to employ SCL decoding with a large list size at high signal-to-noise ratios (SNR) to collect low-weight codewords. Subsequently, [19] refined this method by optimizing memory usage. In parallel, probabilistic approaches with polynomial complexity were introduced in [20, 21] to approximate the weight spectrum of polar codes to strike a balance between accuracy and computational efficiency. 11
@@
To improve the weight spectrum of polar codes, pre-transformed polar codes [22], including CRC-Aided (CA) polar codes[2], parity-check (PC) polar codes[23], and polarization-adjusted convolutional (PAC) codes[24], were proposed. It has been proved that such upper-triangular pre-transformations do not decrease the code distance [22]. The number of minimum-weight codewords for pre-transformed polar codes was calculated in [25]. Building on this foundation, [26] proposed a tree intersection method to reduce the computational complexity. Furthermore, the authors in [27] presented a simplified algorithm to calculate the exact weight spectrum of both original and specific pre-transformed polar codes using coset analysis with exponential complexity. The method was extended to punctured and shortened pre-transformed polar codes in [28]. The average weight spectrum of CA-polar codes, based on the input-output weight enumerating function of original polar codes, was researched in [29]. However, determining the input-output weight enumerating function for original polar codes remains challenging. In [30, 31], the authors proposed efficient recursive formulas to calculate the average weight spectrum of pre-transformed polar codes with polynomial complexity.
@@
Moreover, a method for computing the weight spectrum of rate-compatible polar codes was proposed in [32]. The authors introduced a technique to enumerate all codewords up to a certain Hamming weight for general binary linear block codes. The complexity of using this method is determined by the number of low-weight codewords in the component codes. Although the complexity for polar codes is lower than that of general linear codes, it is still super-polynomial in the code length.
@@
In this paper, we compute the number of minimum-weight codewords for QUP, Wang-Liu shortened, and bit-reversal shortened decreasing polar codes. We also provide iterative formulas for calculating the average weight spectrum of random pre-transformed rate-compatible polar codes. The algorithms operate with polynomial complexity in the code length. This computational characteristic ensures that our methods remain scalable and practical for use with polar codes of varying lengths.
@@ II-A Polar codes as monomial codes
Let $f=x_{1}^{a_{1}}...x_{m}^{a_{m}}$ be a monomial in $\mathcal{M}$, the degree of $f$ is defined as $\text{deg}(f)$, i.e., the Hamming weight of $(a_{1},...,a_{m})$. In particular, 1 is a monomial with degree zero. A monomial $f$ is a factor of a monomial $g$ if $g$ can be written as the product of $f$ and another monomial $h$, i.e., $g=fh$.
@@
and the degree of the polynomial $p=\sum_{e\in\mathcal{M}}u_{f}f$ is defined as $\text{deg}(p)=\max_{u_{f}\neq 0}\{\text{deg}(f)\}$. A polynomial $p$ is said to be linear if deg(p)=1(p)=1.
@@
where the element $p(\bm{a})$ is placed at the $D(\bm{a})$-thth position in $\text{ev}(p)$.
Denote $f_{i}=x_{1}^{a_{1}}...x_{m}^{a_{m}}$ with $i=D(a_{1},...,a_{m})$. Then $\text{ev}(f_{i})$ is exactly the $i$-thth row of $\bm{F}_{N}$. Therefore, for a codeword $\bm{c}=\bm{u}\bm{F}_{N}\in C(\mathcal{I})$, we have $\bm{c}=\text{ev}(p)$ where $p=\sum_{i=1}^{N}u_{i}f_{i}$. In this paper, we sometimes use $p$ to represent codewords $\bm{c}=\text{ev}(p)$. The Hamming weight of $p$ is $\text{wt}(p)\triangleq\text{wt}(\bm{c})$, and $p_{[a,b]}\triangleq\bm{c}_{a}^{b}=(c_{a},c_{a+1},\dots,c_{b})$.
@@
If the maximum degree of monomials in $\mathcal{I}$ is $r$, we call the monomial code $C(\mathcal{I})$ is $r$-thth order.
@@
If the information set is $\mathcal{I}=\{1,x_{1},x_{2},x_{3}\}$, then the generator matrix of the polar code $C(\mathcal{I})$ is
@@
The polynomial $p=x_{3}+x_{1}+1$ is a codeword in $(C(\mathcal{I}))$, which represents $\text{ev}(p)=(1,0,1,0,1,0,0,1)$.
@@ II-E Automorphism group of decreasing monomial codes
The automorphism group Aut(C)(C) is the subgroup of $S_{N}$ containing all the automorphisms of $C$.
@@
A matrix $\bm{A}$ is lower-triangular if $A_{i,i}=1$ and $A_{i,j}=0$ for all $j>i$. The group LTA(m)(m) consists of all the affine transformations $({\bm{A}},{\bm{b}})$ where $\bm{A}$ is lower-triangular. Similarly, a matrix $\bm{A}$ is upper-triangular if $A_{i,i}=1$ and $A_{i,j}=0$ for all $j<i$.
[…]
@@
In [15], it is proved that LTA(m)(m) belongs to the automorphism group of a decreasing monomial code. Later, a larger BLTA group is known to be the affine automorphism group of any specific decreasing monomial code in [36, 35].
@@
When $m=3$, let $\bm{A}=\bm{I}_{3}$ be the identity matrix and $\bm{b}=(1,1,1)$. Then $(\bm{A},\bm{b})$ transforms $x_{i}$ to $x_{i}+1$, which means it interchanges the $i=D((x_{1},x_{2},x_{3}))$-thth bit and the $(N+1-i)=D((x_{1}+1,x_{2}+1,x_{3}+1))$-thth bit for all $1\leq i\leq N$. The corresponding permutation is $(8,7,6,5,4,3,2,1)$.
@@ II-F Puncturing and shortening
Denote $N]=\{1,2,\cdots,N\}$, the difference set of two sets $X$ and $Y$ is denoted as $X/Y$, i.e., $X/Y=\{i\in X\mid i\notin Y\}$. For a codeword $\bm{c}\in\mathbb{F}_{2}^{N}$ and set $X\subseteq[N]$, $\bm{c}_{X}$ is the codeword $\bm{c}$ restricted on $X$, i.e., $\bm{c}_{X}=(c_{i})_{i\in X}$.
@@
Denote the quasi-uniform (QU) sequence to be $\bm{q}=(1,2,\dots,N)$. The QUP pattern $X_{i}$ [4] is the set of the first $i$ bits in $\bm{q}$, i.e., $\{1,2,\dots,i\}$. The QUP polar code with length NiN-i and information set $\mathcal{I}$ is denoted by $C_{P}(\mathcal{I},X_{i})$. The Wang-Liu shortening pattern $Y_{i}$ [5] is the set of the last $i$ bits in $\bm{q}$, i.e., $\{N-i+1,N-i+2,\dots,N\}$. The Wang-Liu shortened polar code with length NiN-i and information set $\mathcal{I}$ is denoted as $C_{S}(\mathcal{I},Y_{i})$.
Denote the bit-reversal sequence $\bm{q}^{\prime}$ to be the bit reversal of $\bm{q}$, i.e., $\bm{q}^{\prime}_{D(a_{m},\dots,a_{1})}=\bm{q}_{D(a_{1},\dots,a_{m})}=D(a_{1},\dots,a_{m})$. The bit-reversal shortening pattern $Y^{\prime}_{i}$ is the last $i$ bits of $\bm{q}^{\prime}$, and the bit-reversal shortened polar code with length NiN-i and information set $\mathcal{I}$ is denoted as $C_{S}(\mathcal{I},Y_{i}^{\prime})$.
@@
Note that when the puncturing pattern $X$ complies with binary domination [37], the symmetric capacity of the monomial $x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ with $D(a_{1},\dots,a_{m})\in X$ becomes zero. When the shortening pattern $Y$ complies with binary domination, the coefficient of monomial $x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ with $D(a_{1},\dots,a_{m})\in Y$ is frozen to zero. Note that QUP, Wang-Liu shortening, and bit-reversal shortening patterns all comply with binary domination. Consequently, for a polynomial $p$ in one of these codes, the coefficient of any monomial $x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ in $p$ with the corresponding index $D(a_{1},\dots,a_{m})$ that is in the puncturing pattern $X$ or the shortening pattern $Y$ is zero. Therefore, with a slight abuse of notation, when the $D(a_{1},\dots,a_{m})$-thth bit is punctured or shortened, we also say the monomial $x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ is punctured or shortened.
In this paper, we focus on the minimum-weight codewords of decreasing QUP, Wang-Liu shortened, and bit-reversal shortened polar codes. We say a shortened polar code is decreasing if the union of the information set and the shortened set is decreasing.
@@
In Section III and IV, we calculate the number of minimum-weight codewords for QUP, Wang-Liu shortened, and bit-reversal shortened polar codes. We provide an outline of the proof in this subsection. Define $T_{P}(f,X)=\{p_{[N]/X}\mid p\in T(f)\}$, which is the set of punctured codewords in $T(f)$ with the puncturing pattern $X$. Similarly, define $T_{S}(f,Y)=\{p_{[N]/Y}\mid p\in T(f),p_{Y}=\bm{0}\}$, which is the set of codewords in $T(f)$ with the shortening pattern $Y$. For $f=x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$, we only need to study $T_{P}(f,X)$ (resp. $T_{S}(f,Y)$) when $D(a_{1},\dots,a_{m})\notin X$ (resp. $Y$) due to Remark 1.
In Section III, we calculate the number of minimum-weight codewords for bit-reversal shortened polar codes $C_{S}(\mathcal{I},Y^{\prime}_{i})$. $T_{S}(f,Y^{\prime}_{i})$ is a subset of $T(f)$ in which the coefficients of monomials in $Y^{\prime}_{i}$ are 0. As shortening does not change the codeword weight, the set of minimum-weight codewords in $C_{S}(\mathcal{I},Y^{\prime}_{i})$ is $\bigcup_{f\in\mathcal{I}_{r}}T_{S}(f,Y^{\prime}_{i})$. We demonstrate that $T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})$ is not empty if and only if the $i$-thth shortened monomial is a factor of $f$. In this case, the number of polynomials removed is $|T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})|=\lambda_{f}/2^{r}$. As $|T_{S}(f,\varnothing)|=|T(f)|=\lambda_{f}$, $|T_{S}(f,Y^{\prime}_{i})|$ can be calculated recursively based on the number of shortened bits.
In Section IV, we analyse the number of minimum-weight codewords for QUP and Wang-Liu shortened polar codes. For QUP, define $M_{P}(\mathcal{I},X_{i})$ to be the set of minimum-weight codewords in $C_{P}(\mathcal{I},X_{i})$. We prove that a codeword in $M_{P}(\mathcal{I},X_{i})$ must also be a minimum-weight codeword before puncturing, meaning it belongs to $\bigcup_{f\in\mathcal{I}}T_{P}(f,X_{i})$. Therefore, we compute the weight spectrum of $\bigcup_{f\in\mathcal{I}}T_{P}(f,X_{i})$ iteratively from the weight spectrum of $T_{P}(f^{\prime},X_{j})$ where $f^{\prime}\preccurlyeq f$ and $j<i$. Additionally, the number of minimum-weight codewords in Wang-Liu shortened polar codes can be inferred from that in QUP codes by symmetry.